OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y

OpenStudy (anonymous):

\[y=x-5\\ y=-x-4\]?

OpenStudy (anonymous):

set them equal \[x-5=-x-4\] and solve for \(x\)

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

must me some typo here or i misunderstood the question

OpenStudy (anonymous):

\[x-5=-x-4\\ 2x-5=-4\\ 2x-1\\ x=\frac{1}{2}\]so i must ahve misunderstood none of those answers are correct

OpenStudy (anonymous):

oh

OpenStudy (solomonzelman):

the signs are missing

OpenStudy (anonymous):

ok i am going to guess that these are inequalities

OpenStudy (anonymous):

oh shott your so right lol let me fix that sorry :/

OpenStudy (solomonzelman):

it is ok... (for the second part of the comment that i understood)

OpenStudy (anonymous):

ooops lol i meant shoot hahah its fixed now :)

OpenStudy (solomonzelman):

i thought I am missing some slang, like shut your ..

OpenStudy (solomonzelman):

shut your so, and i was like "???????...?"\

OpenStudy (anonymous):

oh no lol sorry i didnt spell it right haha

OpenStudy (solomonzelman):

:) it is all good

OpenStudy (anonymous):

does the q make sense now?

OpenStudy (solomonzelman):

graph your inequalities in https://www.desmos.com/calculator and your points graph as well. Wont take long

OpenStudy (solomonzelman):

yes now the q makes sense defnintely

OpenStudy (anonymous):

ok its not working for me lol idk how to use it hahah i was typing it in but it wont give me the answer

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/giouieq801

OpenStudy (solomonzelman):

there is the graph of inequalities and the points.

OpenStudy (anonymous):

ik but i dont know how :/ can you do it for me pls :)

OpenStudy (solomonzelman):

I graphed the points and the inequalities for you. Whichever point lays in the region shaded by both inequalities is the solution.

OpenStudy (anonymous):

ok so what was it?

OpenStudy (anonymous):

@David27

OpenStudy (anonymous):

@ali2x2

OpenStudy (anonymous):

@Baby_Bear69

OpenStudy (ali2x2):

D and you should fan me so you dont tag me when im offline ;)

OpenStudy (anonymous):

ok thats what i was thinking and ok i will :) thnx

OpenStudy (baby_bear69):

I was thinking B, shows what i know.. lol I'm rusty

OpenStudy (anonymous):

lol its ok thnx anyways :)

OpenStudy (teddyiswatshecallsme):

I'm here now C: